SubtoNOT:

Explain in detail with at least two sentences how to represent the difference of z1 and z2 geometrically, given that z1= -7i and z2=4+3i. Then, give the difference z1-z2 in rectangular form.

toga:

To represent the difference of z1 and z2 geometrically, you can plot z1 and z2 as points on the complex plane, and then draw a vector from z2 to z1. The difference z1-z2 in rectangular form is obtained by subtracting the real and imaginary parts of z2 from z1, so z1-z2 = (-7i) - (4+3i) = -4 - 10i.